# User Centric Semantic Communications

### v1

@@ Abstract
Current studies on semantic communications mainly focus on efficiently extracting semantic information to save bandwidth between a transmitter and a user. Although significant process has been made, a fundamental design problem in the semantic communications is that the semantic information is extracted based on certain criteria at the transmitter side along, without considering the user’s actual requirements. As a result, the current semantic communications may cause the loss of critical information that is of primary interest to the user. In such cases, the semantic transmission makes no sense to the user since all the received information is irrelevant. To solve this problem, this paper presents a user centric semantic communication system, where the user sends its request for the desired semantic information to the transmitter at the start of each transmission. Then, the transmitter extracts the required semantic information accordingly. A key challenge is how the transmitter can understand the user’s requests for semantic information and extract the required semantic information in a reasonable and robust manner. We solve this challenge by designing a well-structured framework and leveraging off-the-shelf products, such as GPT-4, along with several specialized tools for detection and estimation. Evaluation results demonstrate the feasibility and effectiveness of the proposed user centric semantic communications.
@@ I Introduction
The concept of semantic communications was originally introduced by Warren Weaver in 1953, focusing on how precisely do the transmitted symbols convey the desired meaning. By concentrating on semantic content, the number of transmitted symbols can be significantly reduced, thereby increasing transmission efficiency in wireless communications. The semantic communications examine the alignment between the transmitter’s intended meaning and the user’s interpretation of the received symbols[1]. A key problem left unsolved in [1] is how to select transmitted symbols that best represent the transmitter’s intended meaning while ensuring alignment with the user’s interpretation.
In [2] and its subsequent works [3, 4, 5, 6, 7], machine learning (ML) techniques such as autoencoders[8] and transformers[9] act as key enablers for the semantic communications. Typically, a neural network-based encoder at the transmitter extracts essential features of the input data, where the dimension of the features is significantly lower than that of the input data. The extracted essential features are defined as the semantic information of the input data [2, 3, 4, 5, 6, 7], which represents the transmitter’s intended meaning. A neural network-based decoder at the user side performs interpretation on the received semantic information by reconstructing the original input data from the transmitter based on it. An end-to-end pre-training is performed beforehand to improve the alignment between the transmitter’s intended meaning and the user’s interpretation.
The current semantic communication framework[2, 3, 4, 5, 6, 7] may lead to the loss of critical information for the user. As illustrated in Fig. 1, although the main content of a video frame is preserved at the user side during the semantic communications, critical details for the user such as a license plate number have been lost. This lost information might be the only detail the user intends to extract from the video frame. As a result, this semantic transmission makes no sense to the user since all the received information is irrelevant, as the information the user cares about has been lost. The loss of critical information problem arises partly due to the out-of-distribution challenge inherent in the ML models when the distribution of an unseen task is different from the distribution of the training data. In this case, the current framework inadvertently discarding subtle yet important information to the user. The out-of-distribution challenge frequently occurs in practice since it is impossible to train an ML model for every possible data distribution.
@@
Figure 1: An autoencoder with a compression rate of 0.5 used in semantic communications fails to recover the license plate number of interest to the user due to the out-of-distribution problem. The peak signal-to-noise ratio (PSNR) of (b) is 15. 97 dB. The autoencoder is pre-trained on the MNIST dataset[10] which consists of 70000 handwritten digits images.
@@
The above observation illustrates a fundamental design flaw in the current semantic communication framework. That is, the current framework extracts the so-called semantic information (i.e., compresses input data) based on certain criteria at the transmitter side along, such as PSNR in image transmissions, and does not take the user’s actual requirement into account. If the transmitter were aware of the user’s information needs in advance, it could take actions to prevent the loss of the user’s required information.
@@
In this paper, we present a user centric semantic communication system, where fulfilling the user’s information requirements is central to each semantic transmission. In the proposed user centric system, the semantic information is redefined as the specific information that user intends to get through a particular data transmission. For example, through the transmission of the video frame in Fig. 1, the user wants to get the license plate number. Thus, the license plate number is the semantic information in this video frame at this particular transmission. In user centric semantic communications, each transmission is initialized by the user. Specifically, at the start of each transmission, the user sends its request on the semantic information to the transmitter. Then, the transmitter extracts the related semantic information and sends it to the user.
@@
We restrict our discussion on the semantic communication in scenarios where the original data at the transmitter is surveillance videos, as the user typically have clear requests for semantic information on the surveillance videos, such as checking traffic situations on a particular road. To achieve the user centric communication system, there are at least two significant challenges.
@@
The first significant challenge is how the transmitter can understand the user’s requests for semantic information and extract the required semantic information from the surveillance video accordingly. To address this, we leverage a large language model (LLM) to understand the user’s requests (termed as tasks), and perform task planning. Specifically, we build a system which consists of a LLM agent and several specialized tools such as an object detection tool. Based on its understanding of the semantic information request, the LLM determines which specialized tools to use and the order of their execution. Then, guided by the task planning, the specialized tools work together to extract the semantic information. Finally, the required semantic information is sent back to the user in text form.
@@
The second significant challenge is the possibility that the LLM’s task planing ability is not robust such that the generated task plan is not reasonable sometimes, and thus cannot fulfill the user’s request directly. To make the user still get the required semantic information, a simple but inefficient way is to have the transmitter send the original video to the user, which can lead to significant bandwidth waste. To solve this challenge, we propose a task reflection process where the LLM assesses whether the current task plan can meet the user’s request on semantic information. If the task reflection result is negative, we alternatively ask the LLM and the specialized tools to select the most relevant video frames related to the user’s request from the surveillance video, and send them back to the user. Upon receiving the video frames, the user can get the required semantic information from the video frames by themselves.
@@
We implement the user centric semantic communication system using the off-the-shell LLM product GPT-4 released by OpenAI [11], along with several specialized tools that have specific detection or estimation capabilities. A demo video is available at [12]. We invite 18 postgraduates from 5 Universities to evaluate the overall system’s performance through traffic surveillance videos. The results show that our system can fulfill 83.90% of the users’ requests for semantic information from the traffic surveillance videos. In addition, compared with the traditional system where the whole videos are transmitted for information extraction at the user side, our system can reduce the number of transmitted video frames by 81.70%, and video data size by 66.33%.
@@
In summary, the proposed user centric semantic communication system successfully meets users’ requirements for semantic information in most cases while significantly reducing the transmitted data size. Second, the system is independent of the data distributions of the surveillance video, as the LLM and specialized tools are off-the-shelf products, meaning we do not need to train them using our data. Third, while we discuss the user centric semantic system in the context of surveillance videos, the framework and methods have the potential to be applied to scenarios where users have clear requests for semantic information from the multimodal data stored at the transmitter.
@@ II Overview of Framework
In this section, we overview the framework of the proposed user centric semantic communication system as shown in Fig. 2. The user and transmitter are located separately and communicate with each other through a wireless network. We focus specifically on the semantic information transmission in the context of surveillance video, since the user always has clear requirements on the semantic information regarding the surveillance videos. The communication is triggered by the user. Specifically, at the start of each transmission, the user sends its request on the semantic information of the surveillance video to the transmitter. Then, the transmitter extracts the related information from the surveillance video, and sends the information to the user.
Figure 2: Overall framework of the proposed user centric semantic communication system.
@@ II-A Transmitter Side
We begin by introducing the transmitter components in the framework shown in Fig. 2. The surveillance videos, e.g., traffic surveillance videos, that are concerned by the user are stored at the transmitter side. Apart from this, the transmitter is connected with an LLM, and a toolbox consisted with several specialized tools. In this paper, GPT-4, released by OpenAI [11], serves as the LLM. In addition, denote the toolbox by
@@
4 Traffic Sign Detection[15] Detects 45 types of traffic signs, e.g, bicycle path, and pedestrian crossing.
@@
The user’s request could be complex and diverse, making it impossible to fulfill with a single specialized tool. The transmitter should first understand the user’s request, and then breaks down the complex request into several smaller and manageable steps (referred to as task planning ), in which each step is executable by a specialized tool. Executing these steps sequentially ensures effective task completion. The transmitter achieves this by leveraging LLM. Specifically, the transmitter relays the user’s request to the LLM, prompting it to perform task planning for semantic information extraction. A key challenge in task planning is ensuring that the LLM can perform planning in a reasonable and robust manner, so the resulting task plan reliably fulfills the user’s request. We will show the task planning design in Section III-A.
Then, the transmitter performs task reflection to determine whether we use the current task plan to extract semantic information for the user, or not. In the case the transmitter uses a wrong task plan, the user will not get the required semantic information. In this case, alternative actions should be taken promptly, preventing any negative impact on the system’s performance. Therefore, the task reflection plays a crucial role in our system. The design of the task reflection is shown in Section III-B. There are two outcomes from the task reflection:
1) The current task plan can fulfill the user’s request. In this case, task execution is performed step-by-step, generating the required semantic information in text form. The transmitter then sends this text-based information to the user. Please refer to Example 1.
2) The current task plan cannot fulfill the user’s request. In this case, another round of task planning is triggered. In the case all the task plans cannot fulfill the user’s request, to make the user still get the required semantic information, a simple but inefficient way is to have the transmitter send the original video to the user, which can lead to significant bandwidth waste. To address this issue, we propose leveraging the LLM and specialized tools for frame selection, i.e., select the most relevant video frames related to the user’s request from the surveillance video. Through the selected frames, the user can get the required information by themselves. We will show how to achieve this in Section IV. Last, the transmitter sends the related video frames to the user. Please refer to Example 2.
@@
6:         Send the text-based information to the user.
@@
12:Perform frame selection and send the frames to user.
Example 1 (Fulfillment). User’s request: “Is there a traffic jam in the video”? Transmitter’s feedback : “I can fulfill your request on the semantic information directly. Based on the analysis of the video, there is no traffic jam detected. The traffic density remains 0.0 for most of the video. Towards the end, it slightly increases to 0.01, but this is still considered light traffic. Therefore, we can confidently say that there is no traffic jam in the video”.
Example 2 (Non-fulfillment). User’s request: “Did an accident happen in the video”? Transmitter’s feedback : “I cannot fulfill your request on the semantic information directly. However, I send the most relevant video frames to you. You may get the required semantic information from the attached video frames (as shown in Fig. 3)”.
II-B User Side
The user’s role is to send requests for semantic information about the surveillance videos of interest and to interpret the feedback from the transmitter: If the feedback from the transmitter is in text form, the user can obtain the required semantic information directly; If the feedback is in image form, the user can get the required semantic information by interpreting the images. For example, in the context of Example 2, by interpreting the video frames from the Fig. 3, the user can determine that an accident has occurred from the second selected video frame shown in Fig. 3(b).
@@ II-B User Side
Figure 3: Selected video frames where the objects highlighted in red boxes may be of interest to the user.
@@ III-A Task Planning
We first introduce the design in task planning through LLM at the transmitter. Task planning with LLMs involves breaking down complex user requests into smaller and manageable steps. However, relying solely on the LLM to perform task planning without any guidance or structured design is insufficient. Specifically, simply inputting the user’s request into the LLM and expecting it to generate a reasonable plan often fails because LLM has a limited understanding of the user’s specific requirements, which can result in impractical or unfeasible steps within the task plan. Additionally, LLM struggles with maintaining long-term dependencies and complex sequencing of steps, resulting in disjointed or out-of-order actions[17]. For example, the first step of the task plan might be mistakenly placed as the third step, rendering the entire plan unreasonable.
@@
  • Video Sampler: LLM samples the input video frames that are used in the next step, according to the user’s request and video frame information (e.g., frame rate and total number of frames). For example, the user may be interested in the semantic information on a particular time of the video.
  • Tool Selection: LLM selects only one appropriate tool in the toolbox $T$ in (1), according to the user’s request. Then, the tool is executed to perform detection or estimation, according to the understanding of the LLM.
  • Analysis: LLM responds to the user’s request based on the execution results from Video Sampler and Tool Selection
@@
Second, a key step in the above task planning design is enabling the LLM to select one suitable specialized tool from the toolbox $T$ during the Tool Selection step. To facilitate this, we first provide the LLM with the toolbox $T$, allowing it to choose from existing tools rather than generating one. Additionally, to ensure the LLM selects one appropriate tool, we include a detailed description for each tool in the toolbox $T$, which encompasses: 1) the functions the tool can perform; 2) the functions it cannot perform; 3) the output format, along with an example. The detailed description of each tool is shown in Appendix -C. Below is an example of the tool Vehicle Density Estimation.
@@
We implement the above task planning through the prompt engineering technique[18] in LLM, which is summarized in Algorithm 2.111When implementing Line 2 of Algorithm 2, we included a detailed description of the three steps involved in executing the prompt. To save space, we omit them here. Through the prompt, along with the other processes in the framework shown in Fig. 2, the Algorithm 2 can output a task plan for the user’s request. For example, the task plan for the request “Is there a traffic jam in the video?” is shown in Line 4 in Algorithm 2.
@@
1:Input: {User’s Request}, and {Toolbox $T$ with Detailed Description}.
2:You are equipped with a toolbox $T$. To fulfill the user’s request, you need to plan first. The plan consists of following three steps in order: Video Sampler, Tool Selection, and Analysis.
@@
1:Input: {User’s Request}, {Current Task Plan}, and {Toolbox $T$ with Detailed Description}.
@@
3:Do you think the current task plan {Current Task Plan} can exactly address the user’s request? You must answer with “Yes” or “No”. Also, you must provide your explanations on this answer.

### v2

# Receiver-Centric Generative Semantic Communications

@@ Abstract
This paper investigates semantic communications between a transmitter and a receiver, where original data, such as videos of interest to the receiver, is stored at the transmitter. Although significant process has been made in semantic communications, a fundamental design problem is that the semantic information is extracted based on certain criteria at the transmitter alone, without considering the receiver’s specific information needs. As a result, critical information of primary concern to the receiver may be lost. In such cases, the semantic transmission becomes meaningless to the receiver, as all received information is irrelevant to its interests. To solve this problem, this paper presents a receiver-centric generative semantic communication system, where each transmission is initialized by the receiver. Specifically, the receiver first sends its request for the desired semantic information to the transmitter at the start of each transmission. Then, the transmitter extracts the required semantic information accordingly. A key challenge is how the transmitter understands the receiver’s requests for semantic information and extracts the required semantic information in a reasonable and robust manner. We address this challenge by designing a well-structured framework and leveraging off-the-shelf generative AI products, such as GPT-4, along with several specialized tools for detection and estimation. Evaluation results demonstrate the feasibility and effectiveness of the proposed new semantic communication system.
@@ I Introduction
This paper studies semantic communications between a transmitter and a receiver, where original data, such as videos of interest to the receiver, is stored at the transmitter. The concept of semantic communications was originally introduced by Warren Weaver in 1953, focusing on how precisely do the transmitted symbols convey the desired meaning. By concentrating on semantic content that is extracted from the original data, semantic communications can significantly reduce the transmitted data size, thereby increasing transmission efficiency in wireless communications. Semantic communications examine the alignment between the transmitter’s intended meaning and the receiver’s interpretation of the received symbols[1]. A key problem left unsolved in [1] is how to select transmitted symbols that best represent the transmitter’s intended meaning while ensuring alignment with the receiver’s interpretation.
@@
Figure 1: An autoencoder with a compression rate of 0.5 used in semantic communications fails to recover the license plate number of interest to the receiver due to the out-of-distribution problem. The peak signal-to-noise ratio (PSNR) of (b) is 15. 97 dB. The autoencoder is pre-trained on the MNIST dataset[10] which consists of 70000 handwritten digits images.
In [2] and its subsequent works [3, 4, 5, 6, 7], machine learning (ML) techniques such as autoencoders[8] and transformers[9] act as key enablers for the semantic communications. Typically, a neural network-based encoder at the transmitter extracts essential features of the input data, where the dimension of the features is significantly lower than that of the input data. The extracted essential features are defined as the semantic information of the input data [2, 3, 4, 5, 6, 7], which represents the transmitter’s intended meaning. A neural network-based decoder at the receiver side performs interpretation on the received semantic information by reconstructing the original input data from the transmitter based on it. An end-to-end pre-training is performed beforehand to improve the alignment between the transmitter’s intended meaning and the receiver’s interpretation.
However, the current semantic communication framework[2, 3, 4, 5, 6, 7] may inadvertently result in the loss of critical information for the receiver. As illustrated in Fig. 1, although the main content of a video frame is preserved at the receiver side during the semantic communications, critical details for the receiver such as a license plate number have been lost. This lost information might be the only detail the receiver intends to extract from the video frame. As a result, this semantic transmission becomes meaningless to the receiver since all the received information is irrelevant, as the information the receiver cares about has been lost. The loss of critical information problem arises partly due to the out-of-distribution challenge inherent in the ML models when the distribution of an unseen task is different from the distribution of the training data. In this case, the current framework fails to preserve subtle yet important information to the receiver. The out-of-distribution challenge frequently occurs in practice since it is impossible to train an ML model for every possible data distribution.
@@
The above observation highlights a fundamental design flaw in the current semantic communication framework. That is, the current framework extracts the so-called semantic information (i.e., compresses input data) based on certain criteria at the transmitter side alone, such as peak signal-to-noise ratio (PSNR) for image transmission, without accounting for the receiver’s specific information needs. If the transmitter were aware of the receiver’s information needs in advance, it could take actions to prevent the loss of the transmitter’s required information.
@@
In this paper, we present a receiver-centric generative semantic communication system, where fulfilling the receiver’s information requirements is central to each semantic transmission. In the proposed receiver-centric system, the semantic information is redefined as the specific information that receiver explicitly seeks from a particular data transmission. For example, if the receiver’s goal is to identify a license plate number in a video frame, then the license plate number is the semantic information of that video frame for that data transmission. In receiver-centric semantic communications, each transmission is initialized by the receiver. Specifically, at the start of each transmission, the receiver sends its request on the semantic information to the transmitter. Then, the transmitter extracts the related semantic information and sends it to the receiver. Note that, although we assume the receiver can also send signals, we continue to refer to this terminal as the “receiver” because its primary function is to receive the required semantic information.
@@
We restrict our discussion on the semantic communication in scenarios where the original data at the transmitter is surveillance videos, as receivers often have well-defined information needs in such contexts, such as monitoring specific traffic situations. To achieve the receiver-centric communication system, there are at least two significant challenges, as discussed below.
@@
The first significant challenge is how the transmitter understands the receiver’s requests for semantic information and then extracts the required information from the surveillance video accordingly. To address this, we leverage a large language model (LLM) to understand the receiver’s requests (termed as tasks), and perform task planning. Specifically, we build a system which consists of an LLM and several specialized tools such as an object detection tool. Based on its understanding of the semantic information request, the LLM determines which specialized tools to use and the order of their execution. Then, guided by the task planning, the specialized tools work together to extract the required semantic information, which is then sent back to the receiver in text form.
@@
The second significant challenge is the possibility that the LLM’s task planing ability is not robust such that the generated task plans are not reasonable sometimes, and thus cannot fulfill the receiver’s request directly. To make the receiver still get the required semantic information, a simple but inefficient way is to have the transmitter send the original video to the receiver, which can lead to significant bandwidth waste. To solve this challenge, we propose a task reflection process where the LLM assesses whether the current task plan can meet the receiver’s request on semantic information. If the assessment is negative, we alternatively ask the LLM and the specialized tools to select the video frames most relevant to the receiver’s request and transmit only these frames to the receiver. The receiver can then obtain the necessary information from these selected frames.
@@
We implement our receiver-centric generative semantic communication system using the off-the-shelf generative AI products, such as GPT-4 released by OpenAI [11], along with several specialized tools that have specific capabilities such as object detection, and estimation. A demo video is available at [12]. We invite 18 postgraduates from 5 Universities to evaluate the overall system’s performance through traffic surveillance videos. The results show that our system can fulfill 83.90% of the receivers’ requests for semantic information. In addition, compared with the traditional systems that transmit entire videos for receiver-side information extraction, our system reduces the number of transmitted video frames by 81.70%, and video data size by 66.33%.
@@
In summary, the proposed receiver-centric generative semantic communication system successfully meets the receiver’s requirements for semantic information in most cases while significantly reducing the transmitted data size. Second, the system is independent of the data distributions of the surveillance video, as the LLM and specialized tools used in this paper are off-the-shelf products, meaning we do not need to train them using our data. Third, although this study focuses on surveillance video transmissions, the framework and methodology have broader potential applications in scenarios where the receivers have clear information requirements from multi-modal data stored at the transmitter.
@@ II Overview of Framework
This section provides an overview of the proposed receiver-centric generative semantic communication framework, illustrated in Fig. 2. In this setup, the receiver and transmitter are located separately and communicate with each other over a wireless network. At the start of each transmission, the receiver sends its request for semantic information to the transmitter. Then, the transmitter extracts the relevant information from the surveillance video, and sends it back to the receiver.
Figure 2: Overall framework of the proposed receiver-centric generative semantic communication system.
@@ II-A Transmitter Side
We begin by introducing the transmitter’s components in the framework shown in Fig. 2. The transmitter stores the surveillance videos that the receiver is interested in, e.g., traffic surveillance videos. In addition, the transmitter is connected with an LLM and a toolbox consisted with several specialized tools. In our system, GPT-4 released by OpenAI [11], serves as the LLM. In addition, denote the toolbox by
@@
4 Traffic Sign Detection[15] Detects 45 types of traffic signs, e.g, bicycle path and pedestrian crossing.
@@
Given the complexity and variety of receiver’s requests, fulfilling them often requires combining multiple specialized tools in $T$. The transmitter should first understand the receiver’s request, and then breaks down the complex request into several smaller and manageable steps, in which each step is executable by a specialized tool. This process is referred to as task planning, achieved by leveraging LLM. Specifically, the transmitter relays the receiver’s request to the LLM, prompting it to perform task planning for semantic information extraction. A key challenge in task planning is ensuring that the LLM generates task plans in a reasonable and robust manner, so that the resulting task plans can reliably fulfill the receiver’s requests. We will show the task planning design in Section III-A.
After task planning, the transmitter performs task reflection to determine whether the generated task plan can accurately fulfill the receiver’s request. If the transmitter uses a wrong task plan, the receiver will not get the required semantic information. In this case, alternative actions should be taken promptly, preventing any negative impact on the system’s performance. Therefore, the task reflection plays a crucial role in our system. The design of the task reflection is shown in Section III-B. There are two outcomes from the task reflection:
1) The generated task plan can fulfill the receiver’s request. In this case, the transmitter proceeds with task execution, carrying out the plan steps in sequence to generate the required semantic information, which is then sent to the receiver as text form. Please refer to Example 1.
2) The generated task plan cannot fulfill the receiver’s request. In this case, another round of task planning is triggered. If after several attempts, no plan succeeds in fulfilling the receiver’s request, the transmitter resorts to sending the raw video frames that most closely match the receiver’s request, allowing the receiver to interpret them directly. Specifically, we propose leveraging the LLM and specialized tools for frame selection, i.e., select the most relevant surveillance video frames related to the receiver’s request. Please refer to Example 2. We will detail how to achieve this in Section IV.
@@
6:         Send the text-based information to the receiver.
@@
12:Perform frame selection and send the frames to receiver.
Example 1 (Fulfillment). Receiver’s request: “Is there a traffic jam in the video”? Transmitter’s feedback : “I can fulfill your request on the semantic information directly. Based on the analysis of the video, there is no traffic jam detected. The traffic density remains 0.0 for most of the video. Towards the end, it slightly increases to 0.01, but this is still considered light traffic. Therefore, we can confidently say that there is no traffic jam in the video”.
Example 2 (Non-fulfillment). Receiver’s request: “Did an accident happen in the video”? Transmitter’s feedback : “I cannot fulfill your request on the semantic information directly. However, I send the most relevant video frames to you. You may get the required semantic information from the attached video frames (as shown in Fig. 3)”.
II-B Receiver Side
The receiver’s role is to initiate semantic requests for surveillance videos of interest and interpret the transmitter’s feedback. If the feedback is in text form, the receiver can obtain the required semantic information directly; If the feedback is in image form, the receiver can get the required semantic information by interpreting the images. For example, in the context of Example 2, by interpreting the selected video frames in Fig. 3, the receiver can determine that an accident has occurred from the second selected video frame.
@@ II-B Receiver Side
Figure 3: Selected video frames where the objects highlighted in red boxes may be of interest to the receiver.
@@ III-A Task Planning
We first introduce the design in task planning through an LLM at the transmitter. Task planning with the LLM involves breaking down complex receiver requests into smaller and manageable steps. However, relying solely on the LLM to perform task planning without any guidance or structured design is insufficient. Specifically, simply inputting the receiver’s request into the LLM and expecting it to generate a reasonable plan often fails because LLM has a limited understanding of the receiver’s specific requirements, leading to impractical or unfeasible steps within the task plan. Additionally, LLM struggles with maintaining long-term dependencies and complex sequencing of steps, resulting in disjointed or out-of-order actions[17]. For example, the first step of the task plan might be mistakenly placed as the third step, rendering the entire plan unreasonable.
@@
  • Video Sampler: LLM samples the input video frames that are used in the next step, according to the receiver’s request and video frame information (e.g., frame rate and total number of frames). For example, the receiver may be interested in the semantic information on a particular time of the video.
  • Tool Selection: LLM selects only one appropriate tool in the toolbox $T$ in (1), according to the receiver’s request. Then, the tool is executed to perform detection or estimation, according to the understanding of the LLM.
  • Analysis: LLM responds to the receiver’s request based on the execution results from Video Sampler and Tool Selection.
@@
Second, a key step in the above task planning design is enabling the LLM to select one suitable specialized tool from the toolbox $T$ during the Tool Selection step. To facilitate this, we provide the LLM with a structured description of the toolbox $T$, guiding it to select from existing tools rather than attempting to generate a new one. Each tool in the toolbox is accompanied by a detailed description that includes: 1) the functions the tool can perform; 2) the functions the tool cannot perform; 3) the output format, along with an example. The detailed description of each tool is shown in Appendix -C. Below is an example of the tool for Vehicle Density Estimation.
@@
We implement the above task planning through the prompt engineering technique[18] in the LLM, which is summarized in Algorithm 2.111When implementing Line 2 of Algorithm 2, we included a detailed description of the three steps involved in executing the prompt. To save space, we omit them here. Through the prompt, along with the other processes in the framework shown in Fig. 2, the Algorithm 2 can output a task plan for the receiver’s request. For example, the task plan for the request “Is there a traffic jam in the video?” is shown in Line 4 in Algorithm 2.
@@
1:Input: {Receiver’s Request}, and {Toolbox $T$ with Detailed Description}.
2:You are equipped with a toolbox $T$. To fulfill the receiver’s request, you need to plan first. The plan consists of following three steps in order: Video Sampler, Tool Selection, and Analysis.
@@
1:Input: {Receiver’s Request}, {Current Task Plan}, and {Toolbox $T$ with Detailed Description}.
@@
3:Do you think the current task plan {Current Task Plan} can exactly address the receiver’s request? You must answer with “Yes” or “No”. Also, you must provide your explanations on this answer.